# A residual driven multiscale method for Darcy’s flow in perforated domains

### v1

@@ 1 Introduction
Traditional numerical methods often struggle with this problem because of the high computational cost required to resolve fine-scale heterogeneities and complex geometries. To overcome these challenges, researchers have developed various specialized methods. One prominent approach is to use upscaling techniques, which incorporate microscopic information into coarse-scale models without explicitly resolving fine-scale features. Examples include the homogenization method [30, 3], the heterogeneous multiscale method [25, 1], and the multicontinuum homogenization method [17, 37]. These methods offer different strategies for bridging micro- and macro-scales, typically by solving a series of local “cell” problems.
@@
The mixed generalized multiscale finite element method (MGMsFEM) [9, 40] has since been applied to numerous practical problems, including linear poroelasticity [33], Stokes flow [34], multiphase transport [35], and flow in perforated domains [13, 14]. In [7], a velocity elimination technique was combined with trapezoidal integration [36] in the lowest-order Raviart–Thomas (RT0) finite element space to diagonalize the mass matrix. This reformulation allows the original saddle-point problem to be re-expressed with pressure as the only unknown variable. As a result, a systematic approach was proposed to construct multiscale basis functions that focus solely on the pressure variable, applied independently within disjoint coarse blocks [24, 22, 23]. To further accelerate the solution process, efficient preconditioners [19, 20, 42, 43, 41] was developed to fully exploit the problem’s disjoint, parallelizable structure.
@@ 2.1 Problem setup
$\kappa^{-1}\mathbf{u}+\nabla p&=0,&\text{in }\Omega^{\epsilon},\\ \mathrm{div}(\mathbf{u})&=f,&\text{in }\Omega^{\epsilon}.\end{aligned}\end{cases}$ (1)
@@
[…]
@@ 2.2 Fine-grid discretization
$\mathbf{V}_{h}$ =$=\{\mathbf{v}_{h}\in\mathbf{V}\mid\mathbf{v}_{h}(w)=(\alpha_{1}+\beta_{1}x_{1},\alpha_{2}+\beta_{2}x_{2}),\ \alpha_{i},\beta_{i}\in\mathbb{R},~{}w\in\mathcal{T}_{h}\},$
$Q_{h}$ =$=\{q_{h}\in Q\mid q_{h}(w)~{}\text{is constant},~{}w\in\mathcal{T}_{h}\}.$
@@
[…]
@@ 3 Offline basis functions
Figure 3: Illustration of the coarse grid and the fine-grid.
@@
$\lambda_{1}^{i}\leq\lambda_{2}^{i}\leq\cdots.$
@@ 4 Online basis functions
where $\mathbf{v}$ is the velocity field corresponding to $q$ via (7). The norm of the residual operator $r_{i}$ is then defined by
@@
$\int_{K_{i}}\kappa^{-1}\bm{\gamma}_{i}\mathbf{v}_{h}-\int_{K_{i}}\beta_{i}\mathrm{div}(\mathbf{v}_{h})&=0,&\forall\mathbf{v}_{h}\in\mathbf{V}_{h}(K_{i}),\\ \int_{K_{i}}\mathrm{div}(\bm{\gamma}_{i})q_{h}&=\int_{K_{i}}(f-\mathrm{div}(\mathbf{u}_{\textnormal{ms}}^{n-1}))q_{h},\quad&\forall q_{h}\in Q_{h}(K_{i}),\end{aligned}\end{cases}$ (13)
@@ 5 Analysis
$\int_{\Omega^{\epsilon}}\kappa^{-1}\mathbf{u}_{\textnormal{ms}}\cdot\mathbf{v}_{\textnormal{ms}}-\int_{\Omega^{\epsilon}}\mathrm{div}(\mathbf{v}_{\textnormal{ms}})p_{\textnormal{ms}}&=0,&\forall\mathbf{v}_{\textnormal{ms}}\in\mathbf{V}_{\textnormal{ms}},\\ \int_{\Omega^{\epsilon}}\mathrm{div}(\mathbf{v}_{\textnormal{ms}})q_{\textnormal{ms}}&=\int_{\Omega^{\epsilon}}fq_{\textnormal{ms}},&\forall q_{\textnormal{ms}}\in Q_{\textnormal{ms}}.\end{aligned}\end{cases}$ (16)
@@ 5.1 Error analysis of the offline stage
In this subsection, we analyze the convergence of the offline multiscale solution. The error estimate depends on the smallest eigenvalue $\Lambda$ whose associated eigenfunction is not included in the offline space. Our proof structure follows prior works such as [9, 7], but we provide a more refined and rigorous formulation.
@@
We now define an auxiliary local solution $\left(\widehat{\mathbf{u}},\widehat{p}\right)$ satisfying
$\kappa^{-1}\widehat{\mathbf{u}}+\nabla\widehat{p}&=0,&\text{in }K_{i},\\ \mathrm{div}(\widehat{\mathbf{u}})&=\bar{f},&\text{in }K_{i},\end{aligned}\end{cases}$ (18)
@@
Let $\mathbf{u}_{h}\in\mathbf{V}_{h}$ be the fine-grid solution of (4), and let $\widehat{\mathbf{u}}\in\mathbf{V}_{h}$ be the solution of (18). Then,
@@
$\int_{\Omega^{\epsilon}}\kappa^{-1}|\mathbf{u}_{h}-\mathbf{u}_{\textnormal{ms}}|^{2}\leq\frac{1}{\Lambda}\sum_{i=1}^{N_{c}}a_{i}(\widehat{p},\widehat{p})+\kappa_{\min}^{-1}\sum_{i=1}^{N_{c}}\int_{K_{i}}|f-\bar{f}|^{2},$ (26)
@@
$\int_{\Omega^{\epsilon}}\kappa^{-1}(\widehat{\mathbf{u}}-\mathbf{u}_{\textnormal{ms}})\cdot(\mathbf{u}_{h}-\mathbf{u}_{\textnormal{ms}})=\int_{\Omega^{\epsilon}}(\widehat{p}-p_{\textnormal{ms}})\mathrm{div}(\mathbf{u}_{h}-\mathbf{u}_{\textnormal{ms}})=\int_{\Omega^{\epsilon}}(\widehat{p}-\widehat{p}_{\textnormal{ms}})\mathrm{div}(\mathbf{u}_{h}-\mathbf{u}_{\textnormal{ms}}),$
where $\widehat{p}_{\textnormal{ms}}$ is the $Q_{\textnormal{ms}}$-projection of $\widehat{p}$, and $\widehat{\mathbf{u}}_{\textnormal{ms}}$ is the corresponding velocity field recovered via (7). Then,
$\int_{\Omega^{\epsilon}}\kappa^{-1}(\widehat{\mathbf{u}}-\mathbf{u}_{\textnormal{ms}})\cdot(\mathbf{u}_{h}-\mathbf{u}_{\textnormal{ms}})=\int_{\Omega^{\epsilon}}\kappa^{-1}(\widehat{\mathbf{u}}-\widehat{\mathbf{u}}_{\textnormal{ms}})\cdot(\mathbf{u}_{h}-\mathbf{u}_{\textnormal{ms}}).$ (28)
@@
$\int_{K_{i}}\kappa^{-1}|\widehat{\mathbf{u}}-\widehat{\mathbf{u}}_{\textnormal{ms}}|^{2}=a_{i}(\widehat{p}-\widehat{p}_{\textnormal{ms}},\widehat{p}-\widehat{p}_{\textnormal{ms}})\leq\frac{1}{\lambda_{l_{i}+1}^{i}}\,a_{i}(\widehat{p},\widehat{p}).$ (29)
@@ 5.2 Error analysis of the online stage
Let $(\mathbf{u}_{\textnormal{ms}}^{n-1},p_{\textnormal{ms}}^{n-1})$ be the multiscale solution at the $(n-1)$-th iteration, and let $Q_{\textnormal{ms}}^{n-1}$ denote the corresponding pressure multiscale space. Assume that we add one online basis function $\beta_{i}$ constructed by (13), and define the enriched space by $Q_{\textnormal{ms}}^{n}=Q_{\textnormal{ms}}^{n-1}\oplus\{\beta_{i}\}$. The new multiscale solution $(\mathbf{u}_{\textnormal{ms}}^{n},p_{\textnormal{ms}}^{n})$ satisfies
@@ 6 Numerical results
Here $p_{h}$ and $\mathbf{u}_{h}$ denote the fine-scale pressure and velocity solutions, and $p_{\textnormal{ms}}$ and $\mathbf{u}_{\textnormal{ms}}$ are the corresponding multiscale solutions. In our computations, we employ several additional techniques to further reduce the dimensionality of the multiscale solution space in each local region. In the offline stage, for example, we exclude any of the first $l_{i}$ eigenfunctions from the offline space if their corresponding eigenvalues exceed $100{,}000$. Our analysis indicates that convergence is primarily governed by the eigenfunctions associated with the smallest eigenvalues; hence, those basis functions corresponding to very large eigenvalues (above this threshold) can be safely omitted from the offline multiscale space. We denote by $L$ the prescribed number of offline basis functions per coarse block. Consequently, due to this eigenfunction exclusion and the constraints imposed by the local geometry, the actual number of basis functions per coarse block may be less than $L$. In the online stage, we adopt a stopping criterion for local enrichment: if the local residual falls below $0.1\%$, no new online basis function is constructed. This threshold is applied in both the uniform and adaptive enrichment strategies. Such a criterion avoids constructing unnecessary basis functions in regions where the solution is already sufficiently accurate.
@@ 6.1 Effect of offline parameters
In this subsection, we present offline simulation results obtained using varying numbers of offline basis functions and different coarse partitions. Detailed results for the three test cases are provided in Tables 13. As expected from our numerical analysis, increasing the number of offline multiscale basis functions leads to a reduction in both pressure and velocity errors. Refining the coarse mesh (i.e., using a smaller coarse grid size $H$) also yields smaller errors. This is because using a finer coarse grid, while maintaining the same number of basis functions per coarse block, effectively increases the total dimension of the multiscale solution space, thereby improving accuracy. Another important observation is that for different coarse grid configurations, if the dimension of the multiscale space is the same, the corresponding errors are comparable. For example, in Table 1, a coarse grid of size $H$ (multiscale space dimension 54,574) yields a pressure error of $0.86\%$, whereas a refined coarse grid of size $H/2$ (dimension 52,397) gives a similar pressure error of $1.12\%$. Likewise, for multiscale space dimensions 108,772 ($H$) and 104,620 ($H/2$), the pressure errors are $0.33\%$ and $0.42\%$, respectively. The velocity error $e_{\mathbf{u}}$ exhibits the same trend. Similar trends are observed in cases 2 and 3, as shown in Tables 2 and 3.
@@ 6.3 Effect of Offline Basis Number and Adaptive Parameter $\theta$
We first examine how varying the number of offline basis functions affects the solution. Specifically, we consider five initial offline spaces with $L=1,2,\dots,5$ basis functions per coarse block. In each case, we subsequently perform $20-L$ steps of uniform online enrichment. The corresponding results are displayed in Figure 6, alongside the results from purely offline enrichment. As observed, the online enrichment consistently outperforms the offline strategy in terms of accuracy, corroborating the findings presented in the previous subsection. Furthermore, increasing the number of offline basis functions in the initial space improves the final accuracy of the online solution. However, the marginal improvement becomes negligible once three offline basis functions per coarse block are included.
Next, we explore the effect of different values of the adaptive parameter $\theta$. We consider four cases: $\theta=1,0.8,0.6,0.4$. Each test starts with three offline basis functions per coarse block, followed by $5$, $8$, $12$, and $17$ steps of online enrichment, respectively. The results are reported in Figure 7. As expected, uniform enrichment ($\theta=1$) achieves the fastest convergence per iteration. However, adaptive enrichment leads to more accurate solutions for a fixed multiscale space dimension. This behavior is attributed to the selective allocation of online basis functions in regions with larger residuals, effectively reducing global error. This trend is consistently observed across all three media.

### v2

@@ 1 Introduction
Traditional numerical methods often struggle with this problem because of the high computational cost required to resolve fine-scale heterogeneities and complex geometries. To overcome these challenges, researchers have developed various specialized methods [35, 7, 31, 29, 30, 49]. One prominent approach is to use upscaling techniques, which incorporate microscopic information into coarse-scale models without explicitly resolving fine-scale features. Examples include the homogenization method [34, 3], the heterogeneous multiscale method [26, 1], and the multicontinuum homogenization method [18, 42]. These methods offer different strategies for bridging micro- and macro-scales, typically by solving a series of local “cell” problems.
@@
The mixed generalized multiscale finite element method (MGMsFEM) [10, 45] has since been applied to numerous practical problems, including linear poroelasticity [38], Stokes flow [40], multiphase transport [39], and flow in perforated domains [15, 14]. In [8], a velocity elimination technique was combined with trapezoidal integration [41] in the lowest-order Raviart–Thomas (RT0) finite element space to diagonalize the mass matrix. This reformulation allows the original saddle-point problem to be re-expressed with pressure as the only unknown variable. As a result, a systematic approach was proposed to construct multiscale basis functions that focus solely on the pressure variable, applied independently within disjoint coarse blocks [25, 23, 24]. To further accelerate the solution process, efficient preconditioners [20, 21, 46, 47, 48] were developed to fully exploit the problem’s disjoint, parallelizable structure.
@@ 2.1 Problem setup
$\begin{cases}\begin{aligned} \kappa^{-1}\mathbf{u}+\nabla p&=0,&\text{in }\Omega^{\epsilon},\\ \mathrm{div}(\mathbf{u})&=f,&\text{in }\Omega^{\epsilon}.\end{aligned}\end{cases}$ (1)
@@
$\begin{cases}\begin{aligned} \int_{\Omega^{\epsilon}}\kappa^{-1}\mathbf{u}\cdot\mathbf{v}-\int_{\Omega^{\epsilon}}\mathrm{div}(\mathbf{v})p&=0,&\forall\mathbf{v}\in\mathbf{V},\\ \int_{\Omega^{\epsilon}}\mathrm{div}(\mathbf{u})q&=\int_{\Omega^{\epsilon}}fq,&\forall q\in Q.\end{aligned}\end{cases}$ (2)
@@ 2.2 Fine-grid discretization
$\mathbf{V}_{h}$ $=\{\mathbf{v}_{h}\in\mathbf{V}\mid\mathbf{v}_{h}(w)=(\alpha_{1}+\beta_{1}x_{1},\alpha_{2}+\beta_{2}x_{2}),\ \alpha_{i},\beta_{i}\in\mathbb{R},~w\in\mathcal{T}_{h}\},$
$Q_{h}$ $=\{q_{h}\in Q\mid q_{h}(w)~\text{is constant},~w\in\mathcal{T}_{h}\}.$
@@
$\begin{cases}\begin{aligned} \int_{\Omega^{\epsilon}}\kappa^{-1}\mathbf{u}_{h}\cdot\mathbf{v}_{h}-\int_{\Omega^{\epsilon}}\mathrm{div}(\mathbf{v}_{h})p_{h}&=0,&\forall\mathbf{v}_{h}\in\mathbf{V}_{h},\\ \int_{\Omega^{\epsilon}}\mathrm{div}(\mathbf{u}_{h})q_{h}&=\int_{\Omega^{\epsilon}}fq_{h},&\forall q_{h}\in Q_{h}.\end{aligned}\end{cases}$ (4)
@@ 3 Offline basis functions
Figure 3: Illustration of the coarse grid and the fine-grid. For perforated regions, fine-grid cells that lie entirely inside the perforations are excluded from the computational domain, while the remaining cells adjacent to the perforation boundaries are treated with appropriate boundary conditions.
@@
$0=\lambda_{1}^{i}\leq\lambda_{2}^{i}\leq\cdots.$
@@
The first eigenvalue equals zero, corresponding to a constant basis function, indicating that the pressure multiscale space enriches the coarse-grid $P_{0}$ space. Because it is a low-dimensional subspace of the fine-grid space, local mass conservation holds only at the coarse-grid level.
@@ 4 Online basis functions
where $\mathbf{v}$ is the velocity field corresponding to $q$ via (7).
The norm of the residual operator $r_{i}$ is then defined by
@@
Indeed, the velocity-weighted $L^{2}$-norm of $r_{i}$ can be regarded as a discrete analogue of the $H^{-1}$ norm in the continuous setting, where $\|q_{h}\|_{V_{i}}$ corresponds to a discrete $H^{1}$ norm in the pressure space.
@@
$\begin{cases}\begin{aligned} \int_{K_{i}}\kappa^{-1}\bm{\gamma}_{i}\mathbf{v}_{h}-\int_{K_{i}}\beta_{i}\mathrm{div}(\mathbf{v}_{h})&=0,&\forall\mathbf{v}_{h}\in\mathbf{V}_{h}(K_{i}),\\ \int_{K_{i}}\mathrm{div}(\bm{\gamma}_{i})q_{h}&=\int_{K_{i}}(f-\mathrm{div}(\mathbf{u}_{\textnormal{ms}}^{n-1}))q_{h},\quad&\forall q_{h}\in Q_{h}(K_{i}),\end{aligned}\end{cases}$ (13)
@@ 5 Analysis
$\begin{cases}\begin{aligned} \int_{\Omega^{\epsilon}}\kappa^{-1}\mathbf{u}_{\textnormal{ms}}\cdot\mathbf{v}_{\textnormal{ms}}-\int_{\Omega^{\epsilon}}\mathrm{div}(\mathbf{v}_{\textnormal{ms}})p_{\textnormal{ms}}&=0,&\forall\mathbf{v}_{\textnormal{ms}}\in\mathbf{V}_{\textnormal{ms}},\\ \int_{\Omega^{\epsilon}}\mathrm{div}(\mathbf{v}_{\textnormal{ms}})q_{\textnormal{ms}}&=\int_{\Omega^{\epsilon}}fq_{\textnormal{ms}},&\forall q_{\textnormal{ms}}\in Q_{\textnormal{ms}}.\end{aligned}\end{cases}$ (16)
@@ 5.1 Error analysis of the offline stage
In this subsection, we analyze the convergence of the offline multiscale solution. The error estimate depends on the smallest eigenvalue $\Lambda$ whose associated eigenfunction is not included in the offline space, which reflects the spectral-decay principle commonly used in the analysis of GMsFEM. Our proof structure follows the spectral analysis framework developed in [10, 8]. In particular, we employ a pressure-space spectral decomposition and develop the analysis for the mixed formulation of the Darcy problem in perforated domains.
@@
[…]
$\begin{cases}\begin{aligned} \kappa^{-1}\widehat{\mathbf{u}}+\nabla\widehat{p}&=0,&\text{in }K_{i},\\ \mathrm{div}(\widehat{\mathbf{u}})&=\bar{f},&\text{in }K_{i},\end{aligned}\end{cases}$ (18)
@@
Let $\mathbf{u}_{h}\in\mathbf{V}_{h}$ be the fine-grid solution of (4), and let $\widehat{\mathbf{u}}\in\mathbf{V}_{h}$ be the solution of (18). Then, we have
@@
$\int_{\Omega^{\epsilon}}\kappa^{-1}|\mathbf{u}_{h}-\mathbf{u}_{\textnormal{ms}}|^{2}\leq\frac{\kappa_{\min}^{-1}}{\Lambda}\sum_{i=1}^{N_{c}}a_{i}(\widehat{p},\widehat{p})+\kappa_{\min}^{-1}\sum_{i=1}^{N_{c}}\int_{K_{i}}|f-\bar{f}|^{2},$ (26)
@@
$\int_{\Omega^{\epsilon}}\kappa^{-1}(\widehat{\mathbf{u}}-\mathbf{u}_{\textnormal{ms}})\cdot(\mathbf{u}_{h}-\mathbf{u}_{\textnormal{ms}})=\int_{\Omega^{\epsilon}}(\widehat{p}-p_{\textnormal{ms}})\mathrm{div}(\mathbf{u}_{h}-\mathbf{u}_{\textnormal{ms}})=\int_{\Omega^{\epsilon}}(\widehat{p}-\widehat{p}_{\textnormal{ms}})\mathrm{div}(\mathbf{u}_{h}-\mathbf{u}_{\textnormal{ms}}),$ (28)
where $\widehat{p}_{\textnormal{ms}}$ is the $Q_{\textnormal{ms}}$-projection of $\widehat{p}$.
@@
[…]
@@ 5.2 Error analysis of the online stage
Let $(\mathbf{u}_{\textnormal{ms}}^{n-1},p_{\textnormal{ms}}^{n-1})$ be the multiscale solution at the (n1)(n-1)-th iteration, and let $Q_{\textnormal{ms}}^{n-1}$ denote the corresponding pressure multiscale space. Assume that we add one online basis function $\beta_{i}$ constructed by (13), and define the enriched space by $Q_{\textnormal{ms}}^{n}=Q_{\textnormal{ms}}^{n-1}\oplus\{\beta_{i}\}$. The new multiscale solution $(\mathbf{u}_{\textnormal{ms}}^{n},p_{\textnormal{ms}}^{n})$ satisfies
@@ 6 Numerical results
Here $p_{h}$ and $\mathbf{u}_{h}$ denote the fine-scale pressure and velocity solutions, and $p_{\textnormal{ms}}$ and $\mathbf{u}_{\textnormal{ms}}$ are the corresponding multiscale solutions. In our computations, we employ several additional techniques to further reduce the dimensionality of the multiscale solution space in each local region. In the offline stage, for example, we exclude any of the first $l_{i}$ eigenfunctions from the offline space if their corresponding eigenvalues exceed 100,000100{,}000. Our analysis indicates that convergence is primarily governed by the eigenfunctions associated with the smallest eigenvalues; hence, those basis functions corresponding to very large eigenvalues (above this threshold) can be safely omitted from the offline multiscale space. We denote by $L$ the prescribed number of offline basis functions per coarse block. Consequently, due to this eigenfunction exclusion and the constraints imposed by the local geometry, the actual number of basis functions per coarse block may be less than $L$. In the online stage, we adopt a stopping criterion for local enrichment: if the local residual falls below $0.1\%$, no new online basis function is constructed. This threshold is applied in both the uniform and adaptive enrichment strategies. Such a criterion avoids constructing unnecessary basis functions in regions where the solution is already sufficiently accurate.
@@ 6.1 Effect of offline parameters
In this subsection, we present offline simulation results obtained using varying numbers of offline basis functions and different coarse partitions. Detailed results for the three test cases are provided in Tables 13. As expected from our numerical analysis, increasing the number of offline multiscale basis functions leads to a reduction in both pressure and velocity errors. Refining the coarse mesh (i.e., using a smaller coarse grid size $H$) also yields smaller errors. This is because using a finer coarse grid, while maintaining the same number of basis functions per coarse block, effectively increases the total dimension of the multiscale solution space, thereby improving accuracy. Another important observation is that for different coarse grid configurations, if the dimension of the multiscale space is the same, the corresponding errors are comparable. For example, in Table 1, a coarse grid of size $H$ (multiscale space dimension 54,574) yields a pressure error of $0.86\%$, whereas a refined coarse grid of size $H/2$ (dimension 52,397) gives a similar pressure error of $1.17\%$. Likewise, for multiscale space dimensions 108,772 ($H$) and 104,620 ($H/2$), the pressure errors are $0.33\%$ and $0.42\%$, respectively. The velocity error $e_{\mathbf{u}}$ exhibits the same trend. Similar trends are observed in cases 2 and 3, as shown in Tables 2 and 3.
@@
We also report the computational cost for both offline and online enrichments on the coarse grid with mesh size $H$ in Table 4. All computations were performed on a standard laptop (Intel Core i7 CPU, 32 GB RAM), and the reported times correspond to solving the coarse-scale linear systems associated with the multiscale formulations. For the offline enrichment, the reported times correspond to solving the coarse-scale system for different total numbers of multiscale basis functions per coarse block ($L=1,4,7,10$), without any iterative procedure. In other words, the offline computation does not involve repeated coarse-scale solves as the basis functions are fixed. In contrast, the online enrichment procedure involves multiple iterations, with one coarse-scale problem solved at each iteration. In Table 4, the online timings correspond to 3, 6, and 9 iterations, where new online basis functions are sequentially added to enrich the multiscale space. Naturally, this iterative procedure increases the overall computational cost. Nevertheless, the online approach achieves a much faster reduction of the approximation error for a given multiscale space dimension, reflecting its effectiveness in improving solution accuracy relative to the offline-only approach.
Model 1 Model 2 Model 3
Offline Online Offline Online Offline Online
1 0.0056 0.0056 0.0037 0.0037 0.0046 0.0046
4 0.0277 0.0919 0.0131 0.0330 0.0283 0.0537
7 0.0665 0.2614 0.0363 0.1109 0.0551 0.1721
10 0.1636 0.6397 0.0633 0.2661 0.0881 0.3982
Table 4: CPU time (in seconds) for three models with different numbers of enrichment steps. The fine-grid solver times for Models 1–3 are 1.2039 s, 0.6499 s, and 0.7563 s, respectively.
@@ 6.3 Effect of Offline Basis Number and Adaptive Parameter $\theta$
We first examine how varying the number of offline basis functions affects the solution. Specifically, we consider five initial offline spaces with $L=1,2,\dots,5$ basis functions per coarse block. In each case, we subsequently perform 20L20-L steps of uniform online enrichment. The corresponding results are displayed in Figure 6, alongside the results from purely offline enrichment. As observed, the online enrichment consistently outperforms the offline strategy in terms of accuracy, corroborating the findings presented in the previous subsection. Furthermore, increasing the number of offline basis functions in the initial space improves the final accuracy of the online solution. However, the marginal improvement becomes negligible once three offline basis functions per coarse block are included.
Next, we explore the effect of different values of the adaptive parameter $\theta$. We consider four cases: $\theta=1,0.8,0.6,0.4$. Each test starts with three offline basis functions per coarse block, followed by $5$, 88, 1212, and $17$ steps of online enrichment, respectively. The results are reported in Figure 7. As expected, uniform enrichment ($\theta=1$) achieves the fastest convergence per iteration. However, adaptive enrichment leads to more accurate solutions for a fixed multiscale space dimension. This behavior is attributed to the selective allocation of online basis functions in regions with larger residuals, effectively reducing global error. This trend is consistently observed across all three media. It is also observed that varying $\theta$ produces only minor differences in the overall convergence behavior. This phenomenon can be explained by the high efficiency of the online basis functions: once a region with large residuals is enriched, its error is rapidly reduced, leading to similar enrichment patterns in subsequent iterations even for different $\theta$ values. In practice, $\theta$ controls the trade-off between computational cost and adaptivity. Smaller $\theta$ values reduce the number of local problems solved per iteration, while larger values accelerate convergence at the expense of additional cost. A more systematic investigation of the optimal choice of $\theta$ and the tolerance parameter used for enrichment termination will be considered in future work.